Using Appendix 2 in your textbook, at what temperature (in K) will the reaction, 2CO(g) + O2(g) --> 2CO2(g), change in spontaneity?
Answer #1

Let vs this solve skep by steps problem stip 1:- for the eatropy change unte standard }h८ guen AS° reactin. J/mol e 213.8 %3D-क्मड Calculate Now, AG= AH-TAS ০০ A G= 0 -566x103 ৪. ।- क- T=3270